OpenStudy (kuoministers):

Please help :) Find the rational values of a and b with a > 0 by forming 2 simultaneous equations and solving them by substitution. (part d)

OpenStudy (kuoministers):

\[(a + b \sqrt{7})^{2} = \frac{ 37 }{ 4 } + 3\sqrt{7}\]

OpenStudy (kuoministers):

it would be \[a^{2} +2ab \sqrt{7} + 7b ^{2} = \frac{ 37 }{ 4 } + 3\sqrt{7}\]

OpenStudy (kuoministers):

i can get the 2 simultaneous equations but i cant go any further \[a^{2} + 7b ^{2} = \frac{ 37 }{ 4 }\] \[2ab \sqrt{7} = 3\sqrt{7}\]

OpenStudy (tyteen4a03):

Wolfram|Alpha http://www.wolframalpha.com/input/?i=solve+%28a+%2B+b+%5Csqrt%7B7%7D%29%5E%7B2%7D+%3D+%5Cfrac%7B+37+%7D%7B+4+%7D+%2B+3%5Csqrt%7B7%7D+ didn't give me a solution. Not sure why. Is \((a + b \sqrt{7})^{2} = \frac{ 37 }{ 4 } + 3\sqrt{7}\) all you're given?

OpenStudy (kuoministers):

yea im sure thats all im given

OpenStudy (kuoministers):

anyways answers were b = 1 a = \[\frac{ 3 }{ 2 }\]

OpenStudy (shubhamsrg):

@Kuoministers you've done correctly thus far.

OpenStudy (shubhamsrg):

Why dont you square both sides of the 2nd eqn. You'll then get 4a^2 b^2 = 9 after that, let a^2 = x and y^2 = b See if that helps ?

OpenStudy (shubhamsrg):

And yes, thats the right answer.

OpenStudy (tyteen4a03):

Never seen questions like this before, and W|A http://www.wolframalpha.com/input/?i=a%5E%7B2%7D+%2B+7b+%5E%7B2%7D+%3D+%5Cfrac%7B+37+%7D%7B+4+%7D%2C+2ab+%5Csqrt%7B7%7D+%3D+3%5Csqrt%7B7%7D gave me 4 solutions. Can you explain a bit how you got that answer @shubhamsrg ? Thanks.

OpenStudy (shubhamsrg):

Wolfram gave 4 solutions, but we are required to choose only those solutions where a>0 and is rational + on comparison, we have 2 eqns as @Kuoministers rightly formed. so 2 eqns, 2 variables, easy to solve.

OpenStudy (kuoministers):

im sorry but squaring it doesn't seem to help?

OpenStudy (shubhamsrg):

after squaring 2nd eqn, you get 4a^2 b^2 =9 right? Now let a^2 = x and b^2 =y Your 2 eqns will be modified as x + 7y = 37/4 and 4xy = 9 Following ?

OpenStudy (kuoministers):

yes

OpenStudy (shubhamsrg):

Now you have 2 eqns and 2 variables, you should be able to solve now hmm?

OpenStudy (kuoministers):

ok gimme a min ill try

OpenStudy (kuoministers):

i cant get the right answer... idk why... im using substitution method

OpenStudy (kuoministers):

\[x = \frac{ 9 }{ 4y }\]

OpenStudy (kuoministers):

\[\frac{ 9 }{ 4y } + 7y = \frac{ 37 }{ 4 }\]

OpenStudy (shubhamsrg):

Correct so far. Please proceed.

OpenStudy (kuoministers):

\[\frac{ 9 }{ y } + 28y = 37\]

OpenStudy (kuoministers):

\[9 + 28y ^{2} = 37y\]

OpenStudy (shubhamsrg):

Its 37/4 there*

OpenStudy (kuoministers):

i multiplied the entire equation by 4

OpenStudy (kuoministers):

and y

OpenStudy (shubhamsrg):

ohh sorry,my bad :P please continue..

OpenStudy (kuoministers):

\[28y ^{2} - 37y + 9 = 0\]

OpenStudy (shubhamsrg):

yes, correct.

OpenStudy (shubhamsrg):

split 37 as 28+ 9

OpenStudy (kuoministers):

(28y - 9)(y - 1)

OpenStudy (kuoministers):

btw why wont y = \[\frac{ 9 }{ 28 }\]

OpenStudy (shubhamsrg):

y = b^2 b is rational right ? so y cant take value 9/28 since that'll make b irrational.

OpenStudy (kuoministers):

your right thanks so much

OpenStudy (shubhamsrg):

hmm, glad i could help..